OpenStudy (anonymous):

Mark layered 3 pieces of wood to build the base for a lamp. The pieces were 1/4 inch thick, 5/8 inch thick and 3/16 inch thick. How thick was the base for the lamp?

OpenStudy (anonymous):

So I suppose you want to place all these fractions under a common denominator so that you can add them. The largest denominator is 16, so you'll want to make everything over 16. (4/4)*(1/4)=(4/16), (2/2)*(5/8)...

OpenStudy (amistre64):

add them all up: 1/4 + 5/8 + 3/16 = ? You any good at fractions?

OpenStudy (amistre64):

id gander to say 1 inch and 1/16th of an inch more... but thats my georgia math coming out :)

OpenStudy (anonymous):

Oh, I was just a bit confused. Thanks.